Bent Gregorian Technical Observing Plan: 31 Aug 2009

Observer: Dave T (from Tucson)
SW support: Torsten, Michele, Chris, Paul
AGw support: Jesper (from Potsdam)

Off-Sky verification

AGW -verified
  • Verify that guider images display normal [Leibold]
  • Verify that WFS images display normal [Leibold]
  • Confirm that all motors can be homed and positioned [Leibold]
  • Verify AGw housekeeping data [Leibold]

    NOTE Aug 28th, 2009: AGW #3 is now controllable through lbtmu105. AGW #4 seems to have technical issues, see IT #2298
    NOTE2 (Ceranski): Having spoken with Torsten, he can confirm that both the guider and WFS images display normally. 
    Also he can confirm that the motors on AGW #3 can be homed and positioned, and that the housekeeping data is available.
    

MCS -verified?
  • Encoder reset: Robust and repeatable [Ashby]
  • El encoder readings consistent over full range [Ashby]
  • Az encoder readings consistent over full range [Ashby]

   NOTE (Ceranski): I spoke with Dave Ashby about these tests, and he told me he has run them during his time on the mountain
   the week of Aug. 24-28. I sent him an e-mail asking him to confirm that on this wiki page.
   

HBS -verified?
  • Obtain Az torque plot over full range [Ashby]
  • Obtain El torque plot over full range [Ashby]
  • Verify adequate elevation balance [Ashby]
  • Verify oil temperature at pads relative to ambient [Ashby]
  • Verify no HBS PLC errors present [Ashby]

   NOTE (Ceranski): I spoke with Dave Ashby about these tests, and he told me he has run them during his time on the mountain
   the week of Aug. 24-28. I sent him an e-mail asking him to confirm that on this wiki page.
   

ECS -verified
  • Verify latch pins and Observing door operation [Little]
  • Verify vent door operation [Little]
  • Confirm chiller nominal operation [Little]
  • Verify both compressors nominal operation [Little]

   NOTE (Ceranski): On the night of Sat. Aug. 29th I thoroughly exercised both shutter doors. The latch pins and ball detents on 
   both doors are working correctly and automatically. I opened/closed the doors several times just to confirm that they are
   working robustly. The vent doors are also working well, and I have seen no problems when opening and closing them. The
   chillers are also working well and I have noticed no problems in either the LBT building cooling, nor the utility building
   cooling. Both air compressors also seem to be working normally. However, the SX air compressor did experience a brief
   problem where it tripped out and required the attention of Jay and Kevin to fix. See IssueTrak #2303 for details. This
   compressor seems to be working fine for the time being.
   

ERC
  • Confirm all drives operational [Little]
  • Obtain torque plot for full range [Little]

   NOTE (Ceranski): All Enclosure Rotation drives seem to be operational and working fine. I've tested the building rotation
   around the entire axis of travel a few times and experienced no problems whatsoever. I asked John Little what it would
   take to obtain a torque plot for the full range of building rotation and he related to me that it would be obtainable, but I
   don't believe that he had time lastweek to get this done. I'll follow up with him on Monday and see what we can come up
   with.
   

  • ERC Following error duing slew:
followingerror.gif

  • ERC motion profile:
position.gif

  • Torque during slew:
torque.gif

  • ERC following error during track:
tracking.gif

MV -verified
  • Verify SX mirror surface temperature uniformity = 0.1C [Biddick]

   NOTE: This spec has never been met.  How about if we try 0.5C ?
   NOTE2: OK
   NOTE3 (Ceranski): I will be ventilating the mirrors (both SX and DX) on Friday and Saturday nights. I'm not sure if we'll
   reach the .5C uniformity in that time, but I can confirm that mirror ventilation is working correctly, and given enough
   ventilation time, it will reach decent thermal equilibrium with the ambient.
   

ROT -verified?
  • Obtain torque plot for full motion range [Ashby]
  • Verify adequate servo stability [Ashby]
  • Inspect cable chain for possible issues [Ashby]

   NOTE (Ceranski): I spoke with Dave Ashby about these tests, and he told me he has run them during his time on the mountain
   the week of Aug. 24-28. I sent him an e-mail asking him to confirm that on this wiki page.
   

TER -verified
  • Verify system start-up robustness [Grenz]
  • Verify tip/tilt/focus command execution [Grenz]
  • Verify rotator command execution [Grenz]

   NOTE: At first, the tertiary UMAC amp-faulted while trying to initialize, 
   but it was run through the power cycling procedure and then it initialized 
   correctly, including the homing procedure. It was moved to the 
   "bentGregorianFront" position, and it performed the move correctly.
   It was homed again, then moved again to the same position.
   

RSU (M2) -verified
  • Confirm nominal hexapod motion range [Grenz]
  • Verify cover removed [Little]

   NOTE: The secondary initializes properly. It was moved to its limits, then 
   re-initialized successfully. An arbitrary move was also done, then 
   the unit was re-homed.
   NOTE2 (Ceranski): I can confirm that the RSU cover is indeed removed.
   

PMC -verified
  • Verify SX hardware nominal operation (voltages, pressures) [Biddick]
  • Verify mirror raise (SX) [Biddick]
  • Verify support system operation over full elevation range [Biddick]

   NOTE: Item three needs to be verified by a TO.  However, it was demonstrated
   to work today during testing with Butch V.
   NOTE2: Chris reports problems with DX hardpoints (electrical)
   NOTE3: (Bob): This is likely the same issue with the hardpoint motor driver electronics 
   that has been an intermittent problem for months.  It occurs periodically with all the 
   hardpoints in both cells.  It has caused some loss of observing time.
   NOTE4 (Ceranski): I can confirm now that the PMC and PSF systems are working normally at all elevations. I tested the
   collimation of both mirrors at elevations between 90 and 0 degrees (in 15 degree increments) and everything seems to
   be behaving well.
   

IRTC -verified
  • Verify utilities (cooling, power, fibers) [Little]
  • Verify camera peltier cooling [Thompson]
  • Verify single frame and movie operation [Thompson]
  • Remaining issue: update DD reference file (20090901, dthompson).

On-Sky verification

Execute TCS startup checklist every night
  • Complete by sunset - 1hr [TO]

Startup
  • Confirm preset command working [Thompson]
  • Verify XY orientation on sky using IRAF and IDL scripts [Thompson]

Pointing & collimation correction (10 minutes):
  • Use WT10 or ACT pointing stars with IRTC
  • Select target at mid EL (50-70 degrees)
  • Preset in PARALLACTIC/TRACK mode:
    With AGw: pointto starname gs=0 Tel="ACQUIRE" PA=0 PAM-
    With IRTC: pointto starname gs=0 Tel="TRACK" PA=0 PAM-
  • If using IRTC, turn on "movie" mode
  • Have TO run "ptspiral" from lbtmu01 and
    • ^C when the star appears on the chip
    • Adjust IE and CA to center; manually focus if necessary
  • Preset in POSITION/ACTIVE mode on axis (Persson std)
  • Allow system to collimate (WFS<400nm rms)

  • Verify Active Optics convergence and stability [Thompson]
  • Verify guiding performance and stability [Thompson]

MCS
  • Servo stability under wind load [Ashby]

PF preset
  • Authorize as LBC, send preset and verify rotator trajectory

On-sky TCS SW Tests (B30)
  • B32 - Test 1 - Test XFLIP/YFLIP are properly set for Gregorian instruments at runtime, particularly LUCIFER.
  • B31 - Test 1 - Test PresetTelescope and OffsetPointing (in RA/Dec) with LBCB31
  • B31 - Test 2 - Test the "absorb" functionality for RA/Dec coordinates (Please see... Overlaps with B21 - Test 3)
  • B30 - Test 1 - Proper motion computation now done for the supplied target and guide star
  • B30 - Test 2 - Ensure tracking (LFBG rotator) recovers after an instrument has completed a ROTATOR_MODE=IDLE and switches to ROTATOR_MODE=POSITION.
  • B30 - Test 3 - Ensure guide stars far off-axis can be acquired as in pre-Build 30 observations.
  • B24 - Test 1 - Spherical Motion Test
  • B23-Test 12 - Verify new oacontrol package features with the following changes.
  • B23-Test 13 - Verify that the Tertiary Mirror in the OSS recognizes that a focal station is selected
  • B23-Test 16 - Verify that the all-on-source signal is working
  • B21-Test 3 - Evaluate the updatePointingReference functionality for coordinate types DETXY (pointing origin) and RADEC (on-sky)
  • B17-Test 1 - Evaluate the OffsetPointing functionality with move types both relative and absolute and modes DETXY (pointing origin) and RADEC (on-sky)

-- JoarBrynnel - 20 Aug 2009

Topic revision: r14 - 01 Sep 2009, JoarBrynnel
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback